Fermi National Laboratory


Accelerator Update for May 4, 2005 - May 6, 2005

Wednesday May 4
The day shift began with the Tevatron (TeV) in shot setup from the Recycler stash alone, and with the MiniBooNE and Meson MCenter experiments taking beam.

Operations turned off the MiniBooNE beam around 8:07 AM.

At 8:09 AM, Operations reported that RF technicians were turning off a Booster RF station (BRF3) for repairs.

Operations established (a Recycler only) store 4126 at 8:47 AM with an initial luminosity of 78.81E30.

At 9:04 AM, BRF4 tripped off and took a power supply (EAPS) with it. Experts reported that BRF4 had failed. They would take it off line and try to bring up BRF3 to take its place.

Booster techs reported that they had the EAPS back up by 9:39 AM.

Operations established beam to MCenter at 10:52 AM.

Operations reported at 10:55 AM that NuMI can resume running at full intensity.

An Antiproton Source (Pbar) expert reported at 11:13 AM that he had found a small LCW leak on a Debuncher quadrupole magnet (D2Q16). Operators contacted the water experts.

Operations established beam to NuMI at 11:28 AM.

A RF expert reported at 11:47 AM that he was putting BRF4 back into service.

Water techs reported at 1:41 PM that they had repaired the Pbar magnet LCW leaks.

Operations tried to turn Pbar back on at 1:44 PM, but a Debuncher Lambertson power supply (D:ELAM) kept tripping its wall breaker. Operators called an expert who had it repaired by 2:26 PM.

Pbar resumed stacking at 2:39 PM.

Main Injector (MI) experts began a study period at 3:18 PM.

At 4:28 PM, the MI safety system tripped off due to a door interlock. Operators prepared the area for an access to investigate.

At 5:50 PM, the EAPS tripped off for the first of two times in fifteen minutes. It reset okay both times.

MI came back and Pbar resumed stacking at 6:28 PM. NuMI and MCenter were taking beam by 6:30 PM.

Recycler experts began a study period at 7:45 PM.

At 10:19 PM, the Pbar beam permit tripped off for the first of five times in forty minutes due to two power supplies not ramping to their correct values. Operators investigate and reported that the supplies were randomly ramping earlier than they should.

At 10:56 PM, a Linac RF station (LRF3) tripped off the first of three times in seven minutes. It reset okay each time.

At 11:09 PM, a vacuum burst occurred at MI-60. It recovered okay.

Recycler studies ended at 11:52 PM.

Thursday May 5
The midnight shift began with Operations monitoring store 4126, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4126 at 5:30 AM.

Operators began preparing MI, TeV, and other areas for safety system tests at 5:38 AM. Booster was prepared for access.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for access at 5:53 AM.

At 8:30 AM, Operations reported that they suffered a short power glitch. The Linac Klystron RF stations all tripped off. The High Voltage Coordinator reported that ComEd had opened and closed a switch off site, which had caused the glitch.

Operations reported at 10:50 AM that the scheduled D0 power outage was beginning. It lasted about a thirty minutes.

The interlock techs reported at 2:39 PM that they had completed testing for MI, the MI-TeV crossover, TeV F-sector, and the Pbar Prevault area.

CDF came out of access and operators interlocked the area at 3:54 PM.

At 3:46 PM, a NuMI expert reported that their target LCW leak had reopened, but at a low rate. The helium backpressure was keeping the situation stable. Experts will conduct target scans tonight.

The interlock techs reported at 4:10 PM that they had completed all of their testing for today.

Operations reported at 4:34 PM that a MiniBooNE LCW system was still down, which stops them from getting beam.

Operations established beam in the Booster at 5:36 PM.

At 7:57 PM, Operations reported that a vacuum leak was discovered upstream of a MI Lambertson. Vacuum techs began working on repairs.

Vacuum techs reported that they completed their MI vacuum work at 11:21 PM.

Operations established a TeV ramp at 11:24 PM.

CDF experimenters requested an access to the collision hall at 11:35 PM.

Friday May 6
The midnight shift began with Pbar stacking, with the MI vacuum recovering, and with operators interlocking areas.

A NuMI expert decided at 1:23 AM that their target scans would wait until morning.

CDF experimenters came out of access at 1:54 AM.

Operations reported at 2:30 AM that Pbar had resumed stacking, but at a low rate due to the MI vacuum recovery.

At 2:54 AM, Operations reported that stacking stopped due to a MI quadrupole power supply (M:Q109V) problem. Operators called in an expert.

At 2:56 AM, Operations reported that Switchyard (SY) beam was on hold due to a MI septa problem. Operators called in an expert.

A Recycler expert began a study at 3:36 AM.

Two MI experts got the quadrupole power supply back on at 3:40 AM and the MI septa working by 3:50 AM. Pbar stacking resumed.

Operations established beam to MCenter at 4:24 AM.

At 4:47 AM, a Pbar emittance blowup halted stacking. Stacking resumed at 5:14 AM.

The Plan for Friday and the Weekend
The plan for today is to allow the TeV a proton only study-store and then go into a shot setup. The rest of the weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/6/2005   email Fermilab